The force constant of a wire is K and that of another wire of the same material is 2K. when both the wires are stretched by the same force, then the work done is
(a) W2 = W1 (b) W2 = 0.5 W1
(c) W2 = 2W1 (d) W2 = 2W12

Grade:12

1 Answers

Ranjeet Gupta
4 years ago
Work done is stored in the form of energy 
So, P.E = W = ½ Kx2
and here we have to given both the wires are stretched by same force [i.e  stretched by same distance]
Hence, x is constant for both the wires
therefore, W1/W2 = K1/ K2       
W1/W2 = K / 2K              …......[ K1= K , K2 =2K]
ANS.    W2 = 2W1